Two planes leave from chicago at noon. One heads east at 420mph while the other heads west at 413 1/3 mph. At what time will they be 2000 miles apart
Answer by TimothyLamb(4379)

combined speed = 420 + 413.333 = 833.333
---
s = d/t
833.333 = 2000/t
t = 2000/833.333
t = 2.4 hours
t = 2 hours and (60 * 0.4) minutes
t = 2 hours and 24 minutes
---
answer:
at 2:24 PM
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php